Seventy children and 160 adults attended a movie theatre. The total reciepts were $620. One adult ticket and two children tickets cost 7$ Find the price of each type of ticket.
Thanks for your help!

Answer by faceoff57(108)

1A+2C=7
160A+70C=620
A=-2C+7
160(-2C+7)+70C=620
-320C+1120+70C=620
-250C+1120=620
-250C=-500
C=2
CHILDREN TICKETS ARE $2
ADULT TICKETS ARE $3